The density and atomic no. of niobium are 8.57 x 10 kg/m³ and 93, respectively. It has one conduction electron per atom. Calculate the London penetration depth of niobium.
1
Expert's answer
2021-10-28T12:34:45-0400
"\lambda=\sqrt\frac{m_e}{\mu_0ne^2},\\\space\\
\lambda=\sqrt\frac{m_eM}{\mu_0N_A\rho e^2}=22.6\text{ nm}."

M above is the molar mass, 0.093 kg/mol, me is the mass of electron, e is the elementary charge.